OpenStudy (anonymous):

The capacity of a cylinder is 942 mL. The radius of the cylinder is 5 cm. What is the height of the cylinder? 12mL 12cm 60mL 60cm

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

The capacity of a cylinder is: \(\Large V=\pi r^2 h=942 ml=942cm^3\) \(\Large \Rightarrow\) the height of the cylinder: \(\Large h=\frac{V}{\pi r^2} = \frac{942}{5^2 \pi} \approx 12\ (cm) \)